# Identifying the post-pandemic determinants of low performing students in Latin America through interpretable Machine Learning SHAP Values—Insights from PISA 2022

### v1

@@ 4.4 SHAP values—interpretable ML estimates
Shapley values have their origin in game theory (Shapley, 1953), measuring the average marginal contribution of a player in a cooperative game. In the context of ML, the player is interpreted as a characteristic (or attribute), and the cooperative game becomes the prediction task performed by the model. Through recent developments (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017), Shapley values have been adapted as an explicative framework for ML models. For the estimation of the main determinants of learning poverty in the LAC region, in particular, I follow the SHAP method (Lundberg and Lee, 2017). Plainly, the SHAP method locally explains a model output by calculating the degree to which each feature (or variable) contributes to a given prediction value. The prediction for the $i$th sample is:
$\widehat{y}_{i}=\widehat{y}_{\text{base}}+\widehat{y}_{i1}+\widehat{y}_{i2}+...+\widehat{y}_{iJ}$ (5)
and where $\widehat{y}_{\text{base}}$ is the average of predictions across all $i$ samples, $\widehat{y}_{ij}$ is the SHAP value of the $j$ feature ($j$ = 1,…,J; with $J$ being total number of feature) in the $i$th sample measuring how much feature $j$ contribute to the prediction of sample $i$. Positive SHAP values ($\widehat{y}_{ij}>0$) indicates that the given feature $j$ has a positive impact on the predicted value of sample $i$, and vice versa.
@@ 4.5 Analytical steps
The steps for the analysis for the PISA 2022 data are displayed in Figure 1. First —and after carefully processing the PISA data and generating each dataset for the dependent variables— the dataset is divided into a trained dataset (80%) and a test dataset (20%). This split, coupled with the stratified $K$-fold (number of folds = 5) cross-validation, ensures that each fold retains proportionality for $y$ educational outcomes and covariates avoiding skewed results. Second, using the train dataset, I run the hyperparameters search fitting different configurations for the five ML classification models, and obtaining their best ones based on the six evaluation metrics. This yields the stacking model based on the train sample. Third, I re-estimate the chosen models excluding a model if it is weight during stacking is equal to zero (interaction 2 in Figure 1) obtaining the SHAP values. Fourth, the final stacked train data model SHAP estimation is then applied to the test data leading to main features covariates selection globally. In the final step, I produce a local version of SHAP values by examining the profile of leading explanatory variables (covariates//features) for students at the extremes of the distribution of the aggregated SHAP values, i.e., $\sum_{j=1}^{J}\phi^{(i)}_{j}$ equal to the minimum and maximum across all $i$ samples (observations).
@@ 5 Results
This section presents the findings via the SHAP analysis for the two samples at the bottom of the learning distribution: level 0 (no achievement) versus level 1, and level 1 versus level 2, both globally (across all samples) and locally (bottom and high performers students within each working sample), dividing the explanatory variables into two sets: students and family covariates, and school covariates. Before, I assess the five binary classification models through evaluation metrics. Countries estimates are also shown as local explanations for specific students based on the SHAP contributions index.
@@ 5.1 Models’ comparison
I present a comprehensive comparison of the efficacy of the five models using a range of evaluation metrics of Section 4.2 for the two working samples. I use a stratified KK-5 fold method approach with 80% of the PISA data working sample belonging to the trained data and the remaining 20% to the test data. The metrics for all models’ configurations are shown in Tables D1 and D2 (Appendix D). For the level0-level1 comparison dataset (the bottom two performing groups), less parsimonious models outperform other simpler models in terms of the degree of accuracy of the classification of students into achieving levels. Notably, GB and RF models consistently show being more effective in terms of the SP and AUC (columns 5 and 6, Table D1), either using more trees and depth and higher learning rates. Results for the level1-level2 comparison dataset (low performing group’s comparison; Table D2), follow an almost identical pattern, with the GB model in particular showing metrics as a leading model.
Table 4 presents evaluation metrics values for the selected better performing models (columns 1 to 6) and each model’s weight (column 7) for the stacked (ensemble) modelling approach where, in the first interaction, it includes all learners//models. Vitally, the stacking model has the benefit of combining the strengths of the individual models it comprises. Figure 2 displays the AUC for all models, including the stacking model where one can see how the classification and identification of the binary categories of educational outcomes improve by the amalgamation of models.
@@
In the first interaction, Table 4 shows that models (learners) with zero weight —i.e., not contributing to the overall prediction of the stacking model— are logit (for the level0-level1 comparison, Panel A, column 7), and logit and neural net (for the level1-level2 comparison; Panel B, column 7). These models are therefore excluded in the second interaction. This leads to the following two stacking models used for predictions and derivation of Shapley values: lassologit, GB, RF and NN (level0-level1 data), and lassologit, GB and RF (level1-level2 data). Also, note that the weights of the final stacking model suggest the relevance in the overall prediction of GB and lassologit models.
5.2 Whole sample—global explanations. Student, family and school determinants
@@
On the one hand, focusing on the top ten most impactful student//househol explanatory variables with the highest SHAP average values (columns 1 and 3 of Table 5), there is a similarity on determinants for bottom and low performers (with 7 out of 10 covariates being in the top ten for these two groups). Primary repetition, sense of belonging to school and wellbeing support from teachers are key drivers of falling into the lowest achievement group among bottom performers (level 0, no competency, column 1), but not in low performing group comparison (column 3). In the reverse direction, engagement in paid work, the intensity of homework and mother’s educational level are among the top ten SHAP values for level 1 versus level 2 comparison, although not for the bottom two levels of achievement comparison. Furthermore, there are some discrepancies on the SHAP values ranking among the two samples. For instance, family SES and family educational support are the leading top two determinants for low performers; yet, in the case of bottom performers sample, they are ranked fifth and sixth in order of importance. Likewise, the number of digital devices and school climate (ranked second and third) are more relevant for bottom performers, and gender having a larger impact when comparing level 1 versus level 2 achievement. Hence, this suggests distinct factors for those students at risk of falling into either level 0 or level 1 of each sample, and therefore the necessity of different strategies when identifying and targeting those students.
@@
On the other hand, as regards to school characteristics SHAP values (Table 5, columns 2 and 4), school size is equal important for both samples, though the rate of school disadvantage and the STR for pedagogic support are higher in the ranking for low performers in comparison to bottom performers. Notably, school type and government funding are more powerful school features for the low performing sample (column 4) than for the bottom performing sample (column 2), whereas the rate of students with a second language and wider support for inclusivity are more relevant for the latter comparison (level 0 versus level 1). In both groups of samples, policies (ability grouping, transfer, admission) are located at the lower end within the SHAP ranked values.4
Figure 3 displays the distribution of SHAP values for each student and covariate for the bottom performing group using beeswarm (top figures) and heatmap (bottom figures) plots5 which offer a more granular approach than the aggregated SHAP values of Table 5 in terms of impacts on predictions.6 All variables are shown in the order of global feature importance, the first one being the most important and the last being the least important one. Below, I explain some of the main findings emerging from these plots.
@@
Figure 3’s beeswarm plot (on the left) shows how high values (displayed in red) of repetition at primary (=1) have a high positive contribution on the prediction (that is, the probability of student’s achieving level 0 instead of level 1), and an equivalent pattern is found for repetition at lower secondary, speaking other language and for those who started primary late and female students. Most of the number of devices’ SHAP values are low and fairly spread (shown in blue), and they have a positive contribution on the chances of students’ falling into level 0 achievement as well as family SES (poorer students) and any engagement in paid work (mostly blue dots), while for teacher wellbeing support estimates are equally divided between high (red) and low (blue) values, something akin is found too for number of books. This is complemented by the SHAP values heatmap for student//family covariates (bottom left) where one can see that covariates with significant homogenous segments (mostly in red) are repetition, devices number, school climate and safety and minority language spoken at home.
@@
For school’s variables (plots on the right of Figure 3) the beeswarm plot shows that, school’s features specific estimates withholding students moving from level 1 to level 2 (with the majority of SHAP values located in the positive section of the x-axis), are low values of testing methods of students (or no testing), high students’ contextual disadvantage, low values of either the ICT index (or weaker ICT infrastructure) or STR (pedagogic staff). And, in the oppositive direction, high values (shown in red) of autonomy, private school and testing, rate of certified teachers and teachers PD attendance, for instance, can all lead to negative predictions of the model output (or higher chances of reaching level 2 achievement across the three subjects). The heatmap for school covariates, in turn, shows that covariates with darker red (with positive impacts on prediction) or blue (negative impacts on prediction) compact segments are: testing, school size (enrolment) and type, and ICT infrastructure.
@@
5.3 Experiment—local explanations. Students at extremes of SHAP index distribution
@@
In this section, I move away from global explanations, shifting the focus towards local explanations of SHAP values which are given by specific observations. This exercise is very relevant policy-wise. For this, I proceed as follows. First, I obtain the overall average contribution of SHAP values across the sample for each student, namely: $\Phi_{i}=\frac{1}{J}\sum_{j=1}^{J}\phi_{j}^{(i)}$; for $i$ = 1,…,$N$. Secondly, I sort out these values and, finally, I select students with the minimum ($\Phi^{\text{min}}_{i}$) and maximum ($\Phi^{\text{max}}_{i}$) contributions so that to profile the impact of covariates at extremes. Once student pairs are identified, I plot leading covariates impacts for each —and fundamentally— with plots showing the specific value each covariate takes. This exercise is important insofar as profiling learning drivers of students and their values7 and, by choosing students at extremes of the SHAP contributions, it is possible to identify students who are the most likely to be performing at level 0 or level 1 (or at level 1 and level 2 for the low performing sample) and their related covariates pushing students’ performance towards either of them.
5.3.1 Level 0 versus level 1—estimates
@@
Here, I compare the profile of students ($\Phi^{\text{max}}_{i}$ = +7.49 and $\Phi^{\text{min}}_{i}$ = -5.92) for the bottom performing sample. In the top two plots of Figure 4 are SHAP values for the student//family group of covariates. A comparison of Figure 4a and Figure 4b demonstrates how inequality operates at the bottom of the learning distribution. A student with the highest probability of being a level 0 achiever attends a school with bad school climate (index of -0.65 falling into high interval); speak other minority language; has no digital devices; has a very weak sense of belonging to school; repeated in primary; and comes from a very poor family (SES index of -2.33 falls into the bottom half of SES distribution); and is engaged in unpaid work (1 day) and paid work (=2.5; between 2 and 3 days per week) (see: Figure 4a). In the oppositive case, a student whose performance falls into level 1, his//her family is much better off (SES index falling into top half); has books (=3) and high number of devices (=10); is satisfied with life and supported by teachers wellbeing wise; have not repeated a grade and does not work; and comes from smaller families (1 sibling) (Figure 4b).
@@
School covariates, too, are dissimilar for these two students’ profiles (Figure 4c and Figure 4d). Consider the first ten ranked covariates in terms of impacts. Smaller schools (enrolment = 129), with nolangtest variable equals to 1 (hence an Indigenous school) and a huge degree of disadvantage (98%) and with weaker support for inclusivity as well as ICT, no PCs and only a third of teachers certified, are all features of schools linked to level 0 achievement. On the contrary, private schools with low STR for pedologic staff (20.2 staff per student versus 109.5 per student for the comparison), low contextual disadvantage and higher autonomy (index = +1.27) and nearly all PCs connected to the web (90.9%) and nearly 55% more qualified teachers, are all school features pushing up the chances of a bottom performer towards a level 1 achievement.
5.3.2 Level 1 versus level 2—estimates
I repeat the same exercise here for the low performer sample (level 1 and level 2). Results are displayed in Figure 5. First, compared to the student with the highest chance of reaching no more than level 1 (Figures 5a and 5b), the student reaching level 2 with the lowest index ($\Phi^{\text{min}}_{i}$) have larger family educational support (index = +1.11); 4 books (instead of 1) and 12 digital devices (instead of 6); a high index regarding school belonging (+2.37 versus -4.47); not engage in paid work and come from a richer family (SES index = +1.57 versus -0.96); and do twice as much homework (2-3 hours versus less than 1 hour).
@@ 5.3.2 Level 1 versus level 2—estimates
Second, and as previously found for the levels 0 and 1 comparison, the level 1 student with $\Phi^{\text{max}}_{i}$ comes from a small Indigenous rural school (as rate for the nolangtest is equal to 1 and urban = 0), while the student with the highest chance of reaching level 2, comes from a large public school (without language minority students) which is located in an urban (large town) area and with parents involvement fostered by school (index = +0.047) where the latter is not encouraged in the Indigenous school of the level 1 student (index = -0.63) (see Figures 5c and 5d). Level 2 student school’s rates for certification and PD attendance of teachers are high (between 80-95 percent) for the school of student reaching level 2 and they are leading SHAP values, but this is not the case for the other student’s school.

### v2

@@ 4.4 SHAP values–interpretable ML estimates
In the body of educational research using ML (see: Ersozlu et al., 2024; Hilbert et al., 2021), there has been a latest growth in the application of Shapley values to identify main features behind educational performance in learning surveys (e.g., Gómez-Talal et al., 2024; Huang et al., 2024). Importantly, we rely on Shapley values to present leading estimates in Section 5. Shapley values have their origin in game theory (Shapley, 1953), measuring the average marginal contribution of a player in a cooperative game. In the context of ML, the player is interpreted as a characteristic (or attribute), and the cooperative game becomes the prediction task performed by the model. Through recent developments (Štrumbelj and Kononenko, 2014; Lundberg and Lee, 2017), Shapley values have been adapted as an explicative framework for ML models. For the estimation of the main determinants of learning poverty in the LAC region, in particular, we follow the SHAP method (SHapley Additive exPlanation) (Lundberg and Lee, 2017). SHAP values offer a local explanation of a model output by calculating the degree to which each feature contributes to a given prediction value when conditioning on that feature.
Formally, the change on prediction or SHAP value $\phi_{i}$(.) for feature (or covariate) $i$ in model $f$ for data point $x$ is given by:
$\phi_{i}(f,x)=\sum_{R\in\mathcal{R}}\frac{1}{M!}\big(f_{x}(P^{R}_{i}\cup i)-f_{x}(P^{R}_{i})\big)$ (7)
where $\mathcal{R}$ is the set of all feature permutations, $P^{R}_{i}$ is the set of all features before $i$ in the ordering $R$, $M$ is the number of input features, and $f_{x}$ is an estimate of the conditional expectation of the model’s prediction. The additive property of SHAP values means that SHAP values add to the output of the model:
$f(x)=\phi_{0}(f)+\sum_{i=1}^{M}\phi_{i}(f,x)$ (8)
Attributions based on the Shapley interaction index result in a matrix of feature attributions. SHAP values main effects (on the matrix’s diagonal) are obtained as the differences of SHAP values and the off-diagonal elements (i.e., SHAP interactions) for a given feature. That is:
$\Phi_{i,i}=\phi_{i}(f,x)-\underbrace{\sum_{j\neq i}\Phi_{i,j}(f,x)}_{\text{SHAP interactions}}$ (9)
There are various algorithms to define the value function for a set of attributes (e.g., DeepSHAP, LinearSHAP, KernelSHAP, TreeSHAP). In the paper, and due to computational issues and the fact that we use a stacked ML model (which is a combination of classification models), we rely on the LinearSHAP (Chen et al., 2020), the default Explainer of the SHAP Python package.
@@ 4.5 Analytical steps
The steps for the analysis for the PISA 2022 data are displayed in Figure 1. First–and after carefully processing the PISA data and generating a dataset for each dependent variable–the dataset is divided into a trained dataset (80%) and a test dataset (20%). This split, coupled with the stratified $K$-fold (number of folds = 5) cross-validation, ensures that each fold retains proportionality values for $y$ educational outcomes and covariates avoiding skewed results. Second, using the train dataset, we run the hyperparameters search fitting different configurations for the five ML classification models, and we obtain their best model specifications based on evaluation metrics. This yields the stacking model for the train sample. Third, we re-estimate the chosen models excluding a model if its weight during stacking is equal to zero (interaction 2 in Figure 1) obtaining SHAP values. Fourth, the final stacked train data model SHAP estimation is then applied to the test data leading to main features covariates selection at a global level. In the final step, we produce a local version of SHAP values by examining the profile of leading explanatory variables (covariates//features) for students at the extremes of the distribution of the aggregated SHAP values (i.e., $\sum_{j=1}^{J}\phi^{(i)}_{j}$ equal to the minimum and maximum across all $i$ samples or observations).
@@ 5 Results
This section presents the findings via the SHAP analysis for the two samples at the bottom of the learning distribution: level 0 (no achievement) versus level 1, and level 1 versus level 2, both globally (across all samples) and locally (bottom and high performers students within each working sample), dividing the explanatory variables into two sets: (i) students and family covariates, and (ii) school covariates. Before, we evaluate the five binary classification models through evaluation metrics. Countries estimates are also shown as local explanations for specific students based on the SHAP contributions index.
@@ 5.1 Models’ comparison
We present a comprehensive comparison of the efficacy of the five models using a range of evaluation metrics. Metrics for all models’ configurations, the grounds for selection of the best performing version of each of the five ML models, are shown in Appendix C. For the level0-level1 dataset (the bottom two performing groups), less parsimonious models outperform other simpler models in terms of the degree of accuracy of the classification of students into achieving levels. Notably, GB and RF models consistently show being more effective in terms of the SP and AUC (columns 5 and 6, Table C1), by using more trees, depth and higher learning rates. Results for the level1-level2 comparison dataset (low performing groups comparison; Table C2), follow a similar pattern, with the GB model in particular showing metrics as a leading model.
Table 4 displays evaluation metrics estimates for the selected best performing models (columns 1 to 6) as well as each model’s weight (column 7) for the stacked (ensemble) modelling approach where, in the first interaction, it includes all models. Vitally, the stacking model has the benefit of combining the strengths of the individual models it comprises. Figure 2 displays the AUC for all models, including the stacking model where one can see how the classification and identification of the binary categories of educational outcomes improve by the amalgamation of models. In the first interaction, Table 4 shows that models (or learners) with zero weight–i.e., not contributing to the overall prediction of the stacking model–are logit (for the level0-level1 comparison, Panel A, column 7), and logit and NN (for the level1-level2 comparison; Panel B, column 7). These models are therefore excluded in the second interaction. This leads to the following two stacking models used for predictions and derivation of Shapley values: lassologit, GB, RF and NN (level0-level1 data), and lassologit, GB and RF (level1-level2 data). Also, note that the higher estimated weights within the final ensemble models for GB and lassologit models suggest their superior relevance in overall predictions.7
@@
5.2 Whole sample–global explanations. Student, family and school determinants
@@
On the one hand, focusing on the top ten most impactful student and family explanatory variables with the highest SHAP average values (columns 1 and 3 of Table 5), there is a similarity on determinants for bottom and low performers (with 7 out of 10 covariates being in the top ten for these two groups). Primary repetition, sense of belonging to school and wellbeing support from teachers are key drivers of a student chances to fall into the lowest achievement group among bottom performers (level 0, no competency, column 1), but not in low performing group comparison (column 3). In the reverse direction, engagement in paid work, the intensity of homework and mother’s educational level are among the top ten SHAP values for the level 1 and level 2 comparison, although not for the bottom two levels of achievement comparison. Furthermore, there are some discrepancies on SHAP values’ ranking among the two samples. For instance, family SES and family educational support are the top two determinants for low performers; yet, in the case of bottom performers sample, they are ranked fifth and sixth in order of importance. The number of digital devices and school climate (ranked second and third) are also more relevant for bottom performers, and gender has a larger impact when comparing level 1 versus level 2 achievement groups. This suggests distinct factors for those students at risk of falling into either level 0 or level 1 of each sample, and therefore the necessity of differential strategies when identifying and targeting those students.
@@
On the other hand, as regards to school characteristics SHAP values (Table 5, columns 2 and 4), school size is equally important for both samples, although the rate of school disadvantage and the STR for pedagogic support are higher in the ranking for low performers compared to their ranking among bottom performers. Distinctly, school type and government funding are more powerful school features for the low performing sample (column 4) than for the bottom performing sample (column 2), whereas the rate of students with a second language and wider support for inclusivity are more relevant for the latter comparison (level 0 versus level 1). In both groups of samples, policies (ability grouping, transfer, admission) are located at the lower end within the SHAP ranked values.8
Figure 3 displays the distribution of SHAP values for each student and covariate for the bottom performing group using beeswarm (top figures) and heatmap (bottom figures) plots9 which offer a more granular approach than the aggregated SHAP values of Table 5 in terms of impacts on predictions.10 All variables are shown in the order of global feature importance. Below, we outline main findings emerging from these plots.
@@
Figure 3’s beeswarm plot (on the left) shows how high values (displayed in red) of repetition at primary (=1) have a positive contribution on the prediction (that is, the probability of student’s achieving level 0 instead of level 1), and an equivalent pattern is found for repetition at lower secondary, speaking other language and for those who started primary late and female students. Most of the number of devices’ SHAP values are low and fairly spread (shown in blue), and they have a positive contribution on the chances of students’ falling into level 0 achievement as well as family SES (poorer students) and any engagement in paid work (mostly blue dots) while, for teacher’s wellbeing support, estimates are equally divided between high (red) and low (blue) values, something similar is found for number of books. This is complemented by the SHAP values heatmap for student//family covariates (bottom left) where one can see that covariates with significant homogenous segments (mostly in red) are repetition, devices number, school climate and safety and minority language spoken at home.
@@
For school’s variables (plots on the right of Figure 3) the beeswarm plot shows that, school’s features specific estimates withholding students moving from level 1 to level 2 (with the majority of SHAP values located in the positive section of the x-axis), are low values of testing methods of students (or no testing), high students’ contextual disadvantage, low values of either the ICT index (or weaker ICT infrastructure) or STR (pedagogic staff). And, in the oppositive direction, high values (shown in red) are found, for instance, for autonomy, private school and testing, rate of certified teachers and teachers PD attendance.11 The heatmap for school covariates, in turn, shows that covariates with darker red (with positive impacts on prediction) or blue (negative impacts on prediction) compact segments are: testing, school size (enrolment) and school type, and ICT infrastructure.
@@
5.3 Experiment–local explanations. Students at extremes of SHAP index distribution
@@
Here, we move away from global explanations by shifting our focus towards local explanations of SHAP values which are given by specific observations. This exercise is very relevant policy-wise. For this, we proceed as follows. First, we obtain the overall average contribution of SHAP values across the sample for each student, namely: $\Phi_{i}=\frac{1}{J}\sum_{j=1}^{J}\phi_{j}^{(i)}$ (for $i$ = 1,…,$N$) and, secondly, we sort out these values and, finally, we select students with the minimum ($\Phi^{\text{min}}_{i}$) and maximum ($\Phi^{\text{max}}_{i}$) contributions so that to profile the impact of covariates at extremes. Once student pairs are identified we plot leading covariates impacts for each student, with graphs showing specific values that each covariate takes. This exercise is important because, by choosing students at extremes of the SHAP contributions, it is possible to identify students who are the most likely to be performing at level 0 or level 1 (or at level 1 and level 2 for the low performing sample) and their related covariates pushing students’ performance towards either of them.12
5.3.1 Level 0 versus level 1–estimates
@@
We compare the profile of students ($\Phi^{\text{max}}_{i}$ = +7.49 and $\Phi^{\text{min}}_{i}$ = -5.92) for the bottom performing sample. In the top two plots of Figure 4 are SHAP values for the student//family group of covariates. A comparison of Figure 4(a) and Figure 4(b) demonstrates how inequality operates at the bottom of the learning distribution. A student with the highest probability of being a level 0 achiever attends a school with bad school climate (index of -0.65 falling into high interval); speak a minority language; has no digital devices; has a very weak sense of belonging to school; repeated in primary; and comes from a very poor family (SES index of -2.33 falls into the bottom half of SES distribution); and is engaged in unpaid work (1 day) and paid work (=2.5; between 2 and 3 days per week) (see: Figure 4(a)). In the oppositive direction, a student whose performance falls into level 1, comes from a better off family (SES index falling into top half); has books (=3) and high number of devices (=10); is satisfied with life and supported by teachers; has not repeated a grade and does not work; and comes from smaller families (Figure 4(b)).
@@
School covariates are also dissimilar for these two students’ profiles (Figures 4(c) and 4(d)). Consider the first ten ranked covariates in terms of impacts. Smaller schools (enrolment = 129), with nolangtest variable equals to 1 (hence an Indigenous school) and a huge degree of disadvantage (98%) and with weaker support for inclusivity and low ICT investments (no PCs) and only a third of teachers certified, are all features of schools linked to a level 0 achievement student. On the contrary, a student attending a private school with low STR for pedologic staff (20.2 students per staff versus 109.5 for the comparison level 0 student), with low contextual disadvantage and higher autonomy (index = +1.27), nearly all PCs connected to the web (90.9%) and around 55% more qualified teachers, are all school features pushing up the chances of a bottom performer student towards level 1 achievement.
5.3.2 Level 1 versus level 2–estimates
We repeat the same exercise for the low performer sample (level 1 and level 2). Results are displayed in Figure 5. First, compared to a student with the highest chance of reaching no more than level 1 (Figures 5(a) and 5(b)), a student reaching level 2 with the lowest index ($\Phi^{\text{min}}_{i}$) has larger family educational support (index = +1.11); 4 books (instead of 1) and 12 digital devices (instead of 6); a high index regarding school belonging (+2.37 versus -4.47); is not engage in paid work and comes from a richer family (SES index = +1.57 versus -0.96); and do twice as much homework (2-3 hours versus less than 1 hour).
@@ 5.3.2 Level 1 versus level 2–estimates
Second, a level 1 student with $\Phi^{\text{max}}_{i}$ comes from a small Indigenous rural school (as rate for the nolangtest is equal to 1 and urban = 0), while the student with the highest chance of reaching level 2 comes from a large public school (without language minority students) which is located in an urban (large town) area and with parents involvement fostered by school (index = +0.047), where the latter is less encouraged at the Indigenous school where the level 1 student is attending (index = -0.63) (see Figures 5(c) and 5(d)). The school rates for certification and PD attendance of teachers are high (between 80-95 percent) and they are leading SHAP values for the school of the level 2 student, but this is not the case for the comparison level 1 student’s school.